FOURIER TRANSFORM INFRARED SPECTROSCOPY OF THE H_{2}F^{+} ν_{1}, ν_{3} AND ν_{2} BANDS

\underline{R. FUJIMORI}, Y. HIRATA, K. KAWAGUCHI, Department of Chemistry, Faculty of Science, Okayama University, 3-1-1, Tsushima-Naka, Okayama 700-8530, JAPAN; I. MORINO, Satellite Remote Sensing Research Section, Center for Global Environmental Research, National Institute for Environmental Studies Onogawa 16-2, Tsukuba, Ibaraki 305-8506, JAPAN.

Since HF is known as interstellar species, the protonated HF may exist with detectable abundance. Vibration-rotation spectra of H_{2}F^{+} ν_{1} and ν_{3} bands were studied by infrared laser spectroscopy^{a}. The ν_{2} band and pure rotational spectra have not been reported. In this study, we report FTIR spectroscopy of H_{2}F^{+} ν_{1}, ν_{3} and ν_{2} bands. The ion was produced with hollow cathode discharge in F_{2}, He and H_{2} mixture. A simultaneous analysis of FT data combined with laser spectroscopic data was curried out for ν_{1} and ν_{3} bands, to determine ground state molecular constants. Absorption lines in the ν_{2} region were assigned by using the ground state combination differences. Determined molecular constants can be used to predict pure rotational transition frequencies.

aE. Schäfer and R. Saykally, J. Chem. Phys. 81, 15 (1984)